OpenStudy (anonymous):

What is the length of the segment that joins the points: (-2,3) and (-8,-7)? Write answer in simplified radical form.

OpenStudy (anonymous):

2sqrt(17)

OpenStudy (anonymous):

This is how you solve any distance problem. Just realize there's a right triangle there, then use the Pythagorean formula.

OpenStudy (anonymous):

Distance b/w 2 coordinates = \[s \sqrt{(x _{2-}}x _{1)^{2}}+(y _{2-}y _{1})^{2}\] =\[\sqrt{(-8--2)^{2}}+(-7-3)^{2}\] = \[\sqrt{-6^{2}}+-10^{2}\] = \[\sqrt{36+100}\] = \[\sqrt{136}\]